OpenStudy (anonymous):

I have been told that a line is x=6-theta y=theta z=2theta... what does this mean? and how do I find a vector for it? Thanks :)

jimthompson5910 (jim_thompson5910):

Theta is just a parameter. So it's just a helper variable in a way. Let's say theta = 0 that means... x = 6-theta x = 6-0 x = 6 y = theta y = 0 z = 2*theta z = 2*0 z = 0 So we have the ordered triple (6,0,0) when theta = 0

jimthompson5910 (jim_thompson5910):

now let's say theta = 1 x = 6-theta x = 6-1 x = 5 y = theta y = 1 z = 2*theta z = 2*1 z = 2 now we have the point (5,1,2)

jimthompson5910 (jim_thompson5910):

Now what is the vector from (6,0,0) to (5,1,2)?

OpenStudy (anonymous):

so are the points (6,0,0) and (5,12) points on the line? and what is an ordered triple?

jimthompson5910 (jim_thompson5910):

it's like an ordered pair, but for 3d space

OpenStudy (anonymous):

umm...(-1, 1,2)?

jimthompson5910 (jim_thompson5910):

basically coordinates for points in 3d space

OpenStudy (anonymous):

oh ok :)

jimthompson5910 (jim_thompson5910):

I like to write vectors with angle brackets like this <-1, 1,2>

jimthompson5910 (jim_thompson5910):

that vector tells you how to move from point to point it's like the slope, but this is for 3d space

OpenStudy (anonymous):

so the direction vector then?

jimthompson5910 (jim_thompson5910):

so if you were to start at (6,0,0), you can move to (5,1,2) by following the vector

jimthompson5910 (jim_thompson5910):

to define the general equation for this line in 3space, we simply start with the starting point (6,0,0) we can think of this as the vector <6,0,0> because we start at the origin (0,0,0) and point to (6,0,0)....essentially the vector <6,0,0>

jimthompson5910 (jim_thompson5910):

then we add on multiples of the vector <-1,1,2> by tacking on a variable (usually t or s) so we have this equation y = <6,0,0> + t*<-1,1,2>

jimthompson5910 (jim_thompson5910):

sorry I guess y is a bad variable choice, how about q q = <6,0,0> + t*<-1,1,2>

OpenStudy (anonymous):

oh thats right...is that the scalar representation of the line?

jimthompson5910 (jim_thompson5910):

notice this when t = 0 q = <6,0,0> + t*<-1,1,2> q = <6,0,0> + 0*<-1,1,2> q = <6,0,0> + <0,0,0> q = <6,0,0> aka the starting point

jimthompson5910 (jim_thompson5910):

when t = 1 q = <6,0,0> + t*<-1,1,2> q = <6,0,0> + 1*<-1,1,2> q = <6,0,0> + <-1,1,2> q = <6 + (-1), 0+1, 0+2> q = <5,1,2> which is the next point over (if you use integral values of t)

OpenStudy (anonymous):

Thats pretty easy then :)

jimthompson5910 (jim_thompson5910):

yeah it's not too bad

OpenStudy (anonymous):

Thanks :)

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (perl):

hmmm